OpenStudy (anonymous):

Quick Question. I need help understanding steps to an example problem. Medal to most helpful.

OpenStudy (anonymous):

What is the resulting product of (x + (2 + 3i))^2?

OpenStudy (anonymous):

kk. im here

OpenStudy (anonymous):

Let me write it all out.

OpenStudy (anonymous):

x^2-4x+13

OpenStudy (anonymous):

@Carpe_Diem77 @Spinex12

OpenStudy (anonymous):

This is the example problem they gave me. I have to solve a problem similar to it, but I'm stuck because I don't know how they got the third number, 12, in the third step.

OpenStudy (anonymous):

oh......... um.............. i messaged carpe to help she might be coming

OpenStudy (anonymous):

Alright. thanks.

OpenStudy (anonymous):

carpe, do u know?

OpenStudy (anonymous):

ummmm no I would have to take some time to solve it but im a little busy....let @MathHelpPls help you

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

what is ur guess so far

OpenStudy (anonymous):

@StudyBuddy17

OpenStudy (anonymous):

\[[x+(2+3i)]^{2}\] \[x^{2}+2[x \times(2+3i)]+(2+3i)^{2}\] \[x^{2}+2(2x+3x i)+(4+12i+9i)^{2}\] \[x^{2}+4x+6x i+4+12i-9\] \[x^{2}+4x+6x i+12i-5\]

OpenStudy (anonymous):

I don't know where the 12i came from in the third step. I'm thinking...

OpenStudy (anonymous):

@souleaterz

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

@Hero @SolomonZelman @mathmale Could you help me?

OpenStudy (mathmale):

\[[x+(2+3i)]^2\] is essentially the square of a binomial. What is \[(a+b)^2 ?\]

OpenStudy (anonymous):

@dgcfhjmtyrfccvbcfgb should be banned

OpenStudy (anonymous):

i read her other emails

OpenStudy (anonymous):

\[(a + b)^{2} = a^{2} + 2ab + b^{2}\]

OpenStudy (anonymous):

ummm....

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

i couldnt figure it out

OpenStudy (anonymous):

gtg bye

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

(a+b)^2=a^2+2ab+b^2 is correct. Thanks. Following this pattern, expand the following: (x + (2 + 3i))^2 (Treat x as you did 'a' and treat (+3i) as you did 'b.'

OpenStudy (anonymous):

i just don't understand how they got 12i.

OpenStudy (anonymous):

i understand all of the steps except when they inserted 12i in the third step.

OpenStudy (mathmale):

Your \[x^{2}+2[x \times(2+3i)]+(2+3i)^{2}\]

OpenStudy (mathmale):

is fine.

OpenStudy (mathmale):

\[x^{2}+2[x \times(2+3i)]+(2+3i)^{2}\]

OpenStudy (mathmale):

if you multiply out the middle term, you get 2x(2+3i), or 4x+6i. Agree or disagree?

OpenStudy (anonymous):

wouldn't it be 4x + 6xi?

OpenStudy (mathmale):

Note: the last term of your \[x^{2}+2(2x+3x i)+(4+12i+9i)^{2}\]

OpenStudy (mathmale):

is incorrect because you have already squared (2+3i). Please note that the square of this binomial is 4 + 6i +9(-1 See whether this helps you to understand where that 12i comes from.

OpenStudy (mathmale):

I'm sorry to be leaving in mid-stream, but have another commitment to take care of now. Good luck!

OpenStudy (anonymous):

I think I got it. Thanks.